OpenStudy (anonymous):

The birds shown on the graph below are flying toward one another at the same speed and same altitude on a straight line course. The birds start from points A(-4,4) and B(10,-2). How far will each bird fly (to the nearest mile) before they collide, if each grid on the graph represents 5 miles? [1] 8 miles [2] 10 miles [3] 38 miles [4] 40 miles

OpenStudy (shiraz14):

horizontal distance covered = 14*5 = 70 miles vertical distance covered = 6*5 = 30 miles Use Pythagoras theorem to work out the length of the hypotenuse, then divide it by 2 to get the distance travelled by each bird

OpenStudy (anonymous):

i know the answer is 38 but i just dont know how to get that @shiraz14

OpenStudy (shiraz14):

@emilyy123: Do you understand the part about the horizontal & vertical distance in my post above?

OpenStudy (anonymous):

yeah i get that part

OpenStudy (anonymous):

but i dont get the rest

OpenStudy (shiraz14):

Ok, so the distance the birds are travelling is the hypotenuse of a right-angled triangle with a horizontal of 70 miles & vertical of 30 miles. We can show this by an illustration below:  We use Pythagoras Theorem to find the length of the hypotenuse, then divide it by 2 since each bird flies to the mid-point of the hypotenuse before they meet.

OpenStudy (anonymous):

ohh but ur supposed to radical it and then i got 53 which is not the right answer

OpenStudy (anonymous):

you will get a decimal but round it to the nearest whole number and then divide it by 2 because if both birds are flying at the same speed that means they actually meet in the middle of the red line = half of the total miles c